How can you determine Amazon's current financial health using Amazon's  return on assets and price to earning for 2019 as a metric.
Answer #1

Return on assets

- Return on assets (ROA) = Operating Income/ Total assets.

This financial ratio analyse the proportion of profit a company earns in comparision to its total resources. It is commonly defined as net income divided by total assets. Net income considered here is the profit after taxes reflected in Income and expense statement. ROAs over 5% are generally considered good but still it depends upon the industry margin and capital intensity.

Amazon.com's annualized Return on Assets (ROA) for the fiscal year that ended in Dec. 2019 is calculated as:

ROA = Net Income (A: Dec. 2019 ) / ( (Total Assets (A: Dec. 2018 ) + Total Assets (A: Dec. 2019 )) / 2 )
= 11588 / ( (162648 + 225248) / 2 )
= 11588 / 193948
= 5.97 %

Amazon.com's annualized Return on Assets (ROA) for the quarter that ended in Dec. 2019 is calculated as:

ROA = Net Income (Q: Dec. 2019 ) / ( (Total Assets (Q: Sep. 2019 ) + Total Assets (Q: Dec. 2019 )) / 2 )
= 13072 / ( (199099 + 225248) / 2 )
= 13072 / 212173.5
= 6.16 %

Analysis : ROA above 5% is considered good. If we analyse the above ROA of Amazon, percentage is pretty good. This shows Assets invested generate a good return in terms of income.

But only disadvantage here is it considers data for last 12 months because changes and fluctuations in the company's earnings or business cycles drastically affect the ratio drastically. ROA can be affected by events such as by goodwill, split, stock buyback or issuance, and, a company's tax rate and its interest payment. ROA may not reflect the true earning power of the assets.

Price to earnings ratio:

- Price to earnings (P/E) = Share Price/ Earning Per Share (EPS).

This ratio (PE Ratio) shows current investor demand for a company share. It is ratio of the share price relative to the annual net income earned by the firm per share.

A high PE ratio generally indicates increased demand because investors anticipate earnings growth in the future. A P/E ratio of 40 is really high, a P/E ratio of 7 is really low, and a ratio of 14 represents the average over modern history.

Amazon.com has a trailing-twelve-months P/E of 93.57X compared to the Internet - Commerce industry's P/E of 29.03X. So if we consider current P/E ratio, it is very good in comparision to the standard ratio.

Financial Analysis of Amazon for year 2019-

- Revenues for Amazon.com, Inc. rose by 20% to $280.52B for the fiscal year of 2019.

- Cash summary statement for all quaters for 2019:

Title

Dec 31, 2019

Sep 30, 2019

Jun 30, 2019

Mar 31, 2019

Cash From Operating Activities 38,514 18,855 10,963 1,846
Cash From Investing Activities -24,281 -20,745 -15,672 -8,123
Cash From Financing Activities -10,066 -6,495 -4,535 -2,377
Net Change in Cash 4,237 -8,619 -9,208 -8,666

Cash From Operating Activities has increased in the last quater. Whereas figures of cash in investing and financing are negative which means instead of surplus cash, there is shortage of cash. And as we see the shortage of cash is increasing in recent quaters.

Debt Equity statement for all quaters for 2019:

Title

Dec 31, 2019

Sep 30, 2019

Jun 30, 2019

Mar 31, 2019

Total Assets 225,248 199,099 191,351 178,102
Total Liabilities 163,188 142,591 138,290 129,692
Total Equity 62,060 56,508 53,061 48,410

Positive increasing trend is observed in Assets, Liabilities and Equity.